# Pedagogical Alignment of Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable potential in various domains, including education. However, their application in educational contexts requires careful ‘pedagogical alignment’, where these models are tuned to assist students in a step-by-step problem-solving process, rather than providing direct answers. In this paper, we investigate algorithms for learning from human preferences (LHP) to achieve this alignment objective. A key challenge in this process is the lack of preference datasets to guide the alignment. To address this, we propose a novel approach for constructing such a dataset using synthetic data generation techniques. Leveraging this dataset, our experiments reveal the superiority of LHP methods over standard supervised fine-tuning (SFT) for pedagogical alignment. Applying LHP algorithms to Mistral and Zephyr, we observe a significant improvement in their ability to guide students, enhancing the pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT. Additionally, we introduce a unique approach for evaluating the pedagogical effectiveness of LLMs using perplexity analysis. We curate a second preference dataset, comparing the perplexity of LLMs’ responses when offering scaffolded step-by-step guidance versus direct solutions. This allows us to quantify the extent to which LLMs prioritize pedagogically effective actions in their generation. Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences.
@@ 1 Introduction
As Large Language Models (LLMs) become integral to the digital landscape, aligning them with the ‘3H’ principle - honesty, harmlessness, and helpfulness - is crucial [1]. In an educational context, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking [2, 3, 4]. This objective, which we term ‘pedagogical alignment,’ involves decomposing complex problems and providing constructive feedback and hints [5, 6]. Previous adaptation of LLMs for education, such as the CLASS framework [7], relied on supervised fine-tuning (SFT). However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment [8, 9].
@@
In this paper, we address this limitation by modeling pedagogical alignment as a learning from human preferences (LHP) problem, where desired alignment is represented by preferences. A major challenge in this approach is the absence of a preference dataset tailored for the educational context. To overcome this, we devise a method for generating a synthetic preference dataset, extending the conversational dataset generation pipeline used in CLASS [7].
Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) [10], Identity Preference Optimization (IPO) [11], and Kahneman Tversky Optimization (KTO) [12] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [13] and Zephyr [14]. We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses. Our results show that DPO and KTO significantly improve Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
To further investigate the effectiveness of LHP over SFT, we introduce a second preference dataset generation technique. This technique allows us to conduct a perplexity analysis, examining the LLM’s predisposition to provide hints and guidance versus giving direct answers. This analysis is critical in understanding how well the LLM aligns with our definition of pedagogical alignment, which emphasizes providing step-by-step guidance over direct answers. We used this technique to generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity. Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance. While SFT partially corrects this behavior, it doesn’t fully optimize it. However, LHP dramatically optimizes this behavior, promoting guidance over direct solutions. This shift in behavior underscores the effectiveness of LHP in aligning LLMs with educational objectives, providing compelling evidence for its superiority over SFT in fostering effective learning experiences.
@@ 2.1 Algorithms for LHP
Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment. LHP/RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [9] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [15, 16] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
More recently, Direct Preference Optimization (DPO) [10] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model. DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency. DPO is designed to directly leverage a dataset of preferences, represented as ${(x,y_{w},y_{l})}$, where each tuple consists of a prompt $x$, a preferred response $y_{w}$, and a dis-preferred response $y_{l}$. Let, $\pi_{\theta}$ be the LLM being finetuned, and $\pi_{\text{ref}}$ be a reference LLM (generally an SFTed model). Then, DPO’s optimization problem can be written as:
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
@@
However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [11] and Kahneman-Tversky Optimization (KTO) [12].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for cost-effective pedagogical alignment of LLMs.
@@ 2.2 Synthetic Student Data Generation
To create our preference dataset, we adapt the CLASS framework [7] that uses GPT-4 [17] to generate synthetic conversational datasets. These datasets simulate a range of interactions between students and GPT-4 using various prompts. An example prompt is illustrated in Table 1. The prompt instructs the model to first evaluate the student response (Evaluation of Student Response), consider which action to employ (Action based on Evaluation), determine the sub-problem (Subproblem State), and conditioned on these fields generate the reply to the student. While the CLASS framework introduced these fields to enhance interpretability for teachers and other educational stakeholders, we utilize them differently to construct a preference dataset, as we will discuss in the next section.
@@ 3 Preference Data Generation for Pedagogical Alignment
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Eval of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Eval": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
@@
${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$ (5)
@@ 4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and LHP models effectively. The CLASS framework used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train LHP models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college books [18], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and LHP models.
To create the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the LHP training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
@@
Model Metric SFT DPO IPO KTO
Mistral Acc 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 75.3 (73,73,80)
F1 0.34 (0.42,0.24,0.37) 0.56 (0.63,0.37,0.67) 0.48 (0.56,0.33,0.56) 0.55 (0.62,0.37,0.65)
Zephyr Acc 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
F1 0.14 (0.21,0.12,0.10) 0.56 (0.65,0.38,0.66) 0.26 (0.32,0.19,0.28) 0.54 (0.63,0.37,0.63)
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy and average F1 score across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three alignment algorithms - DPO, KTO, and IPO - showed notable gains in accuracy and F1 score across all fields as compared to SFT. Across both models, DPO and KTO exhibited comparable performance, outperforming IPO.
@@ 4.2 Models and and Training Details
We experimented on two different LLMs: Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta). We selected Mistral and Zephyr based on their top performance on diverse NLP benchmarks on OpenLLM leaderboard [19, 20] which evaluate scientific knowledge and education-based question answering capabilities: ARC (reasoning based science questions) [21], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [22], and GSM8K (grade school math word problems) [23]. These benchmarks offer a comprehensive evaluation on the models’ performance, especially in education setting.
Our experiments included training and evaluating SFT and LHP models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + LHP + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [24]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [24]: We employed a learning rate of 2e-5 with the AdamW optimizer [25], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the LHP experiments using TRL [26]. For the choice of LHP hyperparamters, we referred to parameters used to align Zephyr [27]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.

### v3

@@ Abstract
Large Language Models (LLMs), when used in educational settings without pedagogical fine-tuning, often provide immediate answers rather than guiding students through the problem-solving process. This approach falls short of pedagogically best practices and limits their effectiveness as educational tools. We term the objective of training LLMs to emulate effective teaching strategies as ‘pedagogical alignment.’ In this paper, we investigate Learning from Human Preferences (LHP) algorithms to achieve this alignment objective. A key challenge in this process is the scarcity of high-quality preference datasets to guide the alignment. To address this, we propose a novel approach for constructing a large-scale dataset using synthetic data generation techniques, eliminating the need for time-consuming and costly manual annotation. Leveraging this dataset, our experiments with Llama and Mistral models demonstrate that LHP methods outperform standard supervised fine-tuning (SFT), improving pedagogical alignment accuracy by 13.1% and 8.7% respectively. Existing evaluation methods also lack quantitative metrics to adequately measure the pedagogical alignment of LLMs. To address this gap, we propose novel perplexity-based metrics that quantify LLMs’ tendency to provide scaffolded guidance versus direct answers, offering a robust measure of pedagogical alignment. Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences. Code and models are available here.
Equal contribution.
@@ 1 Introduction
Large Language Models (LLMs) are increasingly being aligned with the ‘3H’ principle - honesty, harmlessness, and helpfulness (Askell et al., 2021). However, these definitions often require domain-specific adaptations to be truly effective. In educational contexts, for instance, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking (Paul and Elder, 1995; Prince, 2004; Chi, 2009). This objective, which we term ‘pedagogical alignment,’ involves breaking complex problems into manageable steps and providing hints and scaffolded guidance rather than direct answers (Bransford et al., 2000; Hattie and Timperley, 2007). Previous adaptation of LLMs for education, such as the CLASS framework (Sonkar et al., 2023), relied on supervised fine-tuning (SFT). However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment (Christiano et al., 2017; Ziegler et al., 2019).
@@
In this paper, we propose a novel approach to achieve pedagogical alignment by modeling it as learning from human preferences (LHP). This method allows us to represent desired teaching behaviors as preferences, enabling more nuanced optimization than SFT. However, a significant challenge in applying LHP is the scarcity of high-quality, pedagogically-sound preference data. To address this, we introduce a key innovation: leveraging the structured output from the CLASS framework (Sonkar et al., 2023) to create meaningful preference pairs of pedagogically aligned and misaligned responses. This enables us to generate a large-scale, synthetic preference dataset that transforms abstract tutoring principles into concrete, comparative examples, allowing LHP algorithms to learn effective teaching strategies.
Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) (Rafailov et al., 2023), Identity Preference Optimization (IPO) (Azar et al., 2023), and Kahneman Tversky Optimization (KTO) (Ethayarajh et al., 2024) for pedagogical alignment of three open-source LLMs - Llama Dubey et al. (2024), Mistral (Jiang et al., 2023) and Zephyr (Tunstall et al., 2023). We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses. Our results show that DPO and KTO significantly improve Llama, Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 13.1%, 8.7% and 50.0%, respectively, compared to SFT.
Existing evaluation methods for LLMs in education also fail to adequately quantify pedagogical alignment. To address this gap and compare LHP with SFT, we introduce a second preference dataset generation technique and a novel perplexity-based evaluation approach. This allows us to quantitatively measure an LLM’s tendency to provide step-by-step guidance versus direct answers, offering a robust metric for pedagogical alignment. Using our technique we generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity. Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance. While SFT partially corrects this behavior, it doesn’t fully optimize it. However, LHP dramatically optimizes this behavior, promoting guidance over direct solutions. This shift in behavior underscores the effectiveness of LHP in aligning LLMs with educational objectives, providing compelling evidence for its superiority over SFT in fostering effective learning experiences.
@@ 2.1 Algorithms for LHP
Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment. RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by (Ziegler et al., 2019) introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements (Nakano et al., 2021; Ouyang et al., 2022) have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
More recently, Direct Preference Optimization (DPO) (Rafailov et al., 2023) has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for an explicit reward model. DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency. DPO is designed to directly leverage a dataset of preferences, represented as ${(x,y_{w},y_{l})}$, where each tuple consists of a prompt $x$, a preferred response $y_{w}$, and a dis-preferred response $y_{l}$. Let, $\pi_{\theta}$ be the LLM being finetuned, and $\pi_{\text{ref}}$ be a reference LLM (generally an SFTed model). Then, DPO’s optimization problem can be written as:
$W=\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)},L=\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)},\\ L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[W-L\right]$ (1)
@@
DPO tends to overfit to the preference dataset and is sensitive to hyperparameter tuning. Recently Identity Preference Optimization (IPO) (Azar et al., 2023) was introduced to address DPO’s overfitting issues. With addition of a regularization term to the DPO’s loss function, IPO also trains models to convergence without requiring early stopping using:
$L_{\text{IPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\\ \left[\left(\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2\beta}\right)^{2}\right]$
Both DPO and IPO methodologies necessitate a dataset comprising paired preferences, denoted as ${(x,y_{w},y_{l})}$. The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024) presents a solution to this challenge by introducing a loss function formulated exclusively on the basis of singular instances identified as “good” or “bad”.
All three methods, DPO, IPO, and KTO, require a dataset consisting of prompts, accepted responses, and rejected responses. The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for cost-effective pedagogical alignment of LLMs.
@@ 2.2 Synthetic Student Data Generation
To create our preference dataset, we adapt the CLASS framework (Sonkar et al., 2023) that uses GPT-4 Bubeck et al. (2023) to generate synthetic conversational datasets. These datasets simulate a range of interactions between students and GPT-4 using various prompts. An example prompt is illustrated in Table 1. The prompt instructs the model to first evaluate the student response (Evaluation of Student Response), consider which action to employ (Action based on Evaluation), determine the sub-problem (Subproblem State), and conditioned on these fields generate the reply to the student. While the CLASS framework introduced these fields to enhance interpretability for teachers and other educational stakeholders, we utilize them differently to construct a preference dataset, as we will discuss in the next section.
@@ 3 Preference Data Generation for Pedagogical Alignment
[…]
@@
$\mathbf{y}_{t}^{\text{Rejected}}=\mathbf{y}_{t}^{\mathcal{A}_{S}},\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),$
$\forall t\in\{1,\ldots,N\}\big{\}}$
@@
${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$
@@ 4 Experiments
Model Metric SFT DPO IPO KTO
Llama Acc 64 (62, 61, 70) 77 (74, 74, 84) 74 (71, 70, 81) 75 (73, 71, 82)
F1 0.51 (0.54, 0.44, 0.55) 0.57 (0.62, 0.51, 0.59) 0.54 (0.57, 0.47, 0.57) 0.54 (0.58, 0.47, 0.58)
Mistral Acc 65 (61, 61, 72) 74 (71, 69, 80) 71 (67, 67, 79) 72 (69, 69, 79)
F1 0.47 (0.49, 0.41, 0.5) 0.52 (0.54, 0.47, 0.56) 0.5 (0.53, 0.45, 0.54) 0.51 (0.54, 0.46, 0.53)
Zephyr Acc 23 (22, 21, 26) 73 (71, 70, 79) 72 (70, 69, 78) 75 (72, 72, 81)
F1 0.26 (0.29, 0.23, 0.25) 0.55 (0.58, 0.45, 0.62) 0.52 (0.57, 0.45, 0.55) 0.56 (0.59, 0.47, 0.62)
Table 2: SFT vs DPO, IPO, KTO. Each cell displays the average accuracy and F1 score across three classification fields: Evaluation of Student Response, Action Based on the Evaluation, and Subproblem State. We observe consistent improvements in both accuracy and F1 score for all three alignment algorithms—DPO, KTO, and IPO—compared to their SFT counterparts across the models Llama (Llama-3.1-8B-Instruct), Mistral (Mistral-7B-Instruct-v0.2), and Zephyr (zephyr-7b-beta). Notably, DPO and KTO consistently outperform IPO for all models.
@@ 4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and LHP models effectively. The CLASS framework used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train LHP models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompt to generate the problems, subproblems, and associated hints, which were based on biology topics from OpenStax college textbooks (Clark et al., 2018). These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional $1128$ conversations, resulting in $1738$ conversations. Each conversation, on average, comprises approximately $8$ rounds and $1311$ words. This dataset provides a solid platform for training and evaluating SFT and LHP models.
To partition the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into three partitions: SFT training dataset $D_{c}$ ($600$ conversations), LHP training dataset $D_{p}$ ($600$ conversations), and test dataset $D_{t}$ ($450$ conversations). As each round within the conversation is processed into a separate question-answer (QA) pair, $D_{c}$ contains $4942$ QA pairs, $D_{p}$ contains $4921$ QA pairs, and $D_{t}$ contains $3701$ QA pairs.
@@ 4.2 Models and and Training Details
We experimented on three different LLMs: Llama (Llama-3.1-8B-Instruct), Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta) with beta of 0.1 for DPO, IPO, and KTO across all models. Our experiments included training and evaluating SFT and LHP models. It took around $18$ NVIDIA A6000 GPU hours for one cycle (SFT + Data Generation + LHP + Evaluation). A total of $9$ cycles, which are $162$ GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat (Zheng et al., 2023). For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna (Zheng et al., 2023): We employed a learning rate of $2\times 10^{-5}$ with the AdamW optimizer (Loshchilov and Hutter, 2019), a batch size of $16$, a cosine scheduler, a weight decay of $0.05$, a warmup ratio of $0.1$, and a total of 3 epochs. We conducted the LHP experiments using TRL (von Werra et al., 2020). For the choice of LHP hyperparamters, we use beta of $0.1$, a learning rate of $71e-$ 7 with the AdamW optimizer, a batch size of $16$, a cosine scheduler, a weight decay of $0.05$, a warmup ratio of $0.1$, and a total of $3$ epochs.